# In search of the electron-phonon contribution to total energy

### v1

@@ X Size consistency
It was argued in Ref. Paul2023 that the DFPT formulation of AHC zero-point renormalization, $\Delta\varepsilon_{n\mathbf{k}}^{\rm AHC}$ violates size-consistency, whereby the total energy $E^{\rm Allen,AHC}$ would depend on the unit cell size. To show the size consistency of Eq. (123), we simply need to prove the size consistency of $\Delta\varepsilon_{n{\mathbf{k}}}^{\rm AHC}$. For this, we compute the adiabatic zero-point renormalization using the rigid-ion approximation for the Debye-Waller term [Ponce2014a] and a Fourier interpolation of the perturbed potential [Eiguren2008, Gonze2020]. We evaluate the diagonal [Lihm2020] self-energy at the DFT eigenvalues (on-the-mass shell) and use a Sternheimer expression to replace the sum-over-state expression [Gonze2011, Ponce2015].
[…]
primitive supercell
sym. no sym. sym. no sym.
nsym 48 1 24 1
nkpt 8 64 8 32
Q (eBohr) 2.431104 2.431215 2.431114 2.431116
$\times4$$\times$4 $\mathbf{q}$}-grid (meV)
Active space 4 bands 8 bands
$\Delta\varepsilon^{\rm AHC}$ 114.47 114.38 114.07 114.69
$\Delta\varepsilon^{\rm Fan}$ -1549.54 -1549.64 -1548.93 -1549.53
$\Delta\varepsilon^{\rm DW}$ 1664.01 1664.02 1663.00 1664.22
$\times100$$\times$100 $\mathbf{q}$}-interpolation (meV)
Active space 4 bands 8 bands
$\Delta\varepsilon^{\rm AHC}$ 149.73 150.05 149.24 149.77
$\Delta\varepsilon^{\rm Fan}$ -1577.59 -1577.71 -1577.02 -1577.76
$\Delta\varepsilon^{\rm DW}$ 1727.33 1727.75 1726.27 1727.54
Active space 9 bands 18 bands
$\Delta\varepsilon^{\rm AHC}$ 150.82 151.20 150.36 150.84
$\Delta\varepsilon^{\rm Fan}$ -1578.99 -1579.27 -1578.39 -1579.04
$\Delta\varepsilon^{\rm DW}$ 1729.81 1730.46 1728.75 1729.88
ecut & $\mathbf{k}$ Fan (meV) DW (meV) AHC (meV)
Ha active full full full
symmetry
40 & 43 173.997 -1577.592 1727.326 149.734
60 & 83 176.978 -1582.746 1735.074 152.328
no symmetry
40 & 43 174.305 -1577.707 1727.755 150.048
60 & 83 176.997 -1582.777 1735.082 152.305
[…]
We finish by presenting in Table 4 the values of the ZPR without interpolation of the perturbed potential and conclude that interpolation is not the source of the problem, but in fact symmetry detection. We also remark that exclusively for Table 4, the data were computed with Abinit v9.10.5 and a PBE pseudopotential without NLCC from PseudoDojo [Setten2017], without spin-orbit coupling.
XI Diamond and Lonsdaleite allotropes
@@
Having analyzed and compared the behavior of the theory and implementation in different cases, we now converge the various contributions to the total energies with phonon momentum $\mathbf{q}$}. In this case, we can no longer use FD methods as supercell size quickly becomes prohibitively larger. First, we remark that the various total energy calculations require dense phonon momentum $\mathbf{q}$} grid samplings of the electron-phonon interaction, which can be computationally expansive. Two interpolation methods are often used to accurately and efficiently interpolate $\mathbf{q}$}-dependent properties: the Wannier-Fourier interpolation of the electron-phonon matrix elements [Giustino2007, Ponce2016] and the Fourier interpolation of the perturbed potential [Gonze2020, Brunin2020, Brunin2020a]. Importantly, long-range electrostatics should analytically be removed and added back to ensure accuracy of the interpolation [Gonze1997, Verdi2015, Sjakste2015, Stengel2016, Royo2019, Royo2020, Brunin2020, Brunin2020a, Jhalani2020, Royo2021, Ponce2021, Ponce2023, Ponce2023a].
@@
To analyze the importance of the different terms, we compute these for another allotrope of diamond called lonsdaleite or hexagonal diamond [Bundy1967] which forms naturally under high-pressure and high-temperature and is typically found in meteorite impacts. It crystallizes in the P63/mmc space group with 4 atoms per primitive cell. We computed a DFT-BO crystal structure of a=4.7477 Bohr and c=7.9011 Bohr, slightly above the experimental lattice values of a=4.743 Bohr and c=7.786 Bohr [Bundy1967]. Our computed fractional internal parameter is u=0.374535. The volume per atom is similar for diamond and lonsdaleite, with 38.465 Bohr3/atom and 38.560 Bohr3/atom, respectively. We compute the quadrupole tensor of lonsdaleite integrated on a $\times18$$\times$12 k-point grid and obtain the non-equivalent values: Qκyxx = -1.643 e Bohrs, Qκzxx = -1.663 e Bohrs, Qκxxz = -1.018 e Bohrs, and Qκzzz = 2.775 e Bohrs.
Prior DFT calculations have reported that diamond is more stable than lonsdaleite by 2.6 kcal/mol [Jones2016]. As reported in Table 5, we also find diamond to be more stable but only by -53.964 meV/(2 atoms), which corresponds to 0.6222 kcal/mol. We then study the second (phononic) and fourth-order (electron-phonon only) contributions to the total energy using DFPT and in particular the formulation of Eq. (106) for $E^{\rm elph}$. We report in Fig. 4 the values of $E^{\rm ph}$, $E^{\rm Allen,AHC}$, and $E^{\rm elph}$ as a function of momentum $\mathbf{q}$-grid integration. We also integrate on $\mathbf{k}$-points using crystal symmetries and a $\times8$$\times$8 and $\times6$$\times$4 grids for diamond and lonsdaleite, respectively. We verify that increasing the $\mathbf{k}$-point grids changes the results by less than 1%. Both $E^{\rm ph}$ and $E^{\rm elph}$ converge quickly with $\mathbf{q}$-point grids but $E^{\rm Allen,AHC}$ requires very dense grids. This is not surprising, as this term is obtained by integrating the eigenvalue renormalization, a quantity that is known to converge slowly with phonon momentum grids [Ponce2015, Ponce2025a].
@@
We note that among all the $\lambda^{6}$ order energy terms presented in Table 1, the effect of the $\mathcal{E}^{\rm(6),elm}$ term given in Eq. (162) can be easily computed, approximately, by renormalizing the $E^{\rm ph}$ term as:
@@
In this approximation, we consider the $\Delta m_{\kappa\alpha,\kappa^{\prime}\alpha^{\prime}}^{\textrm{elm},-1}$ tensor to be diagonal, isotropic for each nucleus, with the value needed to fulfill the sum rule mentioned in Eq. (17) of Ref. Scherrer2017, namely, $n_{e\kappa}=6$ for carbon nuclei. We report it in Table 5 and find the very small rounded value of -0.049 meV for both allotropes, which therefore does not affect their relative stability.
@@
Indeed, as already discussed at the end of Sec. IV, the $\{\mathbf{R}^{0}\}$ atomic positions are at their BO minimum, but might no longer be at their minimum once the additional energy terms are included. This is indeed the case as seen in Fig. 5, and the diamond lattice parameter increases from 6.7518 Bohr to 6.7776 Bohr when including the three additional forces, which represent a 0.4% increase. This term is known as zero-point lattice expansion [BrousseauCouture2022]. We emphasize that this is a small but non-negligible increase that occurs at zero Kelvin and will be much larger with the temperature increase (thermal expansion). For londsdaleite, a similar increase would occur, and we do not report it. However, in contrast to diamond, londsdaleite has an internal parameter $u$ and we find in Fig. 5 a reduction of the internal parameter from 0.37459 to 0.37448. Moreover, the lowering of the energy between the optimized atomic position and $\{\mathbf{R}^{0}\}$ is the quasiharmonic energy, defined (for the terms that we address in the present study) as:

### v2

@@ IX Validation for diamond at the phonon zone-center
Finally, it was argued in Ref. 55 that the DFPT formulation of AHC zero-point renormalization, $\Delta\varepsilon_{n\mathbf{k}}^{\rm AHC}$ violates size-consistency, whereby the total energy $E^{\rm Allen,AHC}$ would depend on the unit cell size. To show the size consistency of Eq. (119), we simply need to prove the size consistency of $\Delta\varepsilon_{n{\mathbf{k}}}^{\rm AHC}$. For this, we compute the adiabatic zero-point renormalization using the rigid-ion approximation for the Debye-Waller term [59] and a Fourier interpolation of the perturbed potential [22, 29]. We evaluate the diagonal [43] self-energy at the DFT eigenvalues (on-the-mass shell) and use a Sternheimer expression to replace the sum-over-state expression [27, 60], and report the details in Sec. LABEL:sec:size of the SI. We find excellent agreement, validating size consistency.
X Diamond and Lonsdaleite allotropes
@@
Having analyzed and compared the behavior of the theory and implementation in different cases, we now converge the various contributions to the total energies with phonon momentum $\mathbf{q}$}. In this case, we can no longer use FD methods as supercell size quickly becomes prohibitively larger. First, we remark that the various total energy calculations require dense phonon momentum $\mathbf{q}$} grid samplings of the electron-phonon interaction, which can be computationally expansive. Two interpolation methods are often used to accurately and efficiently interpolate $\mathbf{q}$}-dependent properties: the Wannier-Fourier interpolation of the electron-phonon matrix elements [25, 61] and the Fourier interpolation of the perturbed potential [29, 18, 17]. Importantly, long-range electrostatics should analytically be removed and added back to ensure the accuracy of the interpolation [33, 89, 83, 84, 75, 74, 18, 17, 38, 76, 64, 66, 65].
@@
To analyze the importance of the different terms, we compute these for another allotrope of diamond called lonsdaleite or hexagonal diamond [19] which forms naturally under high-pressure and high-temperature and is typically found in meteorite impacts. It crystallizes in the P63/mmc space group with 4 atoms per primitive cell. We computed a DFT-BO crystal structure of a=4.7477 Bohr and c=7.9011 Bohr, slightly above the experimental lattice values of a=4.743 Bohr and c=7.786 Bohr [19]. Our computed fractional internal parameter is u=0.374535. The volume per atom is similar for diamond and lonsdaleite, with 38.465 Bohr3/atom and 38.560 Bohr3/atom, respectively. We compute the quadrupole tensor of lonsdaleite integrated on a $\times18$$\times$12 k-point grid and obtain the non-equivalent values: Qκyxx = -1.643 e Bohrs, Qκzxx = -1.663 e Bohrs, Qκxxz = -1.018 e Bohrs, and Qκzzz = 2.775 e Bohrs.
Prior DFT calculations have reported that diamond is more stable than lonsdaleite by 2.6 kcal/mol [39]. As reported in Table 3, we also find diamond to be more stable but only by -53.964 meV/(2 atoms), which corresponds to 0.6222 kcal/mol. We then study the second (phononic) and fourth-order (electron-phonon only) contributions to the total energy using DFPT and in particular the formulation of Eq. (102) for $E^{\rm elph}$. We report in Fig. 4 the values of $E^{\rm ph}$, $E^{\rm Allen,AHC}$, and $E^{\rm elph}$ as a function of momentum $\mathbf{q}$-grid integration. We also integrate on $\mathbf{k}$-points using crystal symmetries and a $\times8$$\times$8 and $\times6$$\times$4 grids for diamond and lonsdaleite, respectively. We verify that increasing the $\mathbf{k}$-point grids changes the results by less than 1%. Both $E^{\rm ph}$ and $E^{\rm elph}$ converge quickly with $\mathbf{q}$-point grids but $E^{\rm Allen,AHC}$ requires very dense grids. This is not surprising, as this term is obtained by integrating the eigenvalue renormalization, a quantity that is known to converge slowly with phonon momentum grids [60, 63].
@@
We note that among all the $\lambda^{6}$ order energy terms presented in Table 1, the effect of the $\mathcal{E}^{\rm(6),elm}$ term given in Eq. (LABEL:eq:6el) can be easily computed, approximately, by renormalizing the $E^{\rm ph}$ term as:
@@
In this approximation, we consider the $\Delta m_{\kappa\alpha,\kappa^{\prime}\alpha^{\prime}}^{\textrm{elm},-1}$ tensor to be diagonal, isotropic for each nucleus, with the value needed to fulfill the sum rule mentioned in Eq. (17) of Ref. 81, namely, $n_{e\kappa}=6$ for carbon nuclei. We report it in Table 3 and find the very small rounded value of -0.049 meV for both allotropes, which therefore does not affect their relative stability.
@@
Indeed, as already discussed at the end of Sec. IV, the $\{\mathbf{R}^{0}\}$ atomic positions are at their BO minimum, but might no longer be at their minimum once the additional energy terms are included. This is indeed the case as seen in Fig. 5, and the diamond lattice parameter increases from 6.7518 Bohr to 6.7776 Bohr when including the three additional forces, which represent a 0.4% increase. This term is known as zero-point lattice expansion [16]. We emphasize that this is a small but non-negligible increase that occurs at zero Kelvin and will be much larger with the temperature increase (thermal expansion). For londsdaleite, a similar increase would occur, and we do not report it. However, in contrast to diamond, londsdaleite has an internal parameter $u$ and we find in Fig. 5 a reduction of the internal parameter from 0.37459 to 0.37448. Moreover, the lowering of the energy between the optimized atomic position and $\{\mathbf{R}^{0}\}$ is the quasiharmonic energy, defined (for the terms that we address in the present study) as: